i need help on how to write and equation in slope form using (2,-5) with a slope of -5/3 thanks
Answer by Alan3354(69443)

i need help on how to write and equation in slope form using (2,-5) with a slope of -5/3 thanks
-----------------------
I think you mean thru the point (2,-5) with a slope of -5/3
Use y-y1 = m*(x-x1) where (x1,y1) is the point and m is the slope.
y+5 = (-5/3)*(x-2)
y+5 = (-5/3)x + 10/3
y = (-5/3)x - 5/3 slope-intercept form
-----------------
5x + 3y = -5 standard form